If a force of 50N moves a load of 30kg up an inclined plane. Determine the magnitude of the force if the plane is inclined at an angle of 30
1
Expert's answer
2021-09-14T14:15:49-0400

Let's find the magnitude of the resultant force:


"F_{res}=F_{appl}-mgsin\theta,""F_{res}=50\ N-30\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot sin30^{\circ}=-97\ N."

The sign minus means that the load can't move up an inclined plane.